OpenStudy (anonymous):

H9P1: Response to a Delayed Impulse L=20.0H and C=20.26mF. A=2/π=0.64 Coulombs at time t=9.0s. B. At the initial time what is the total energy, in Joules, stored in the circuit? C. At the time just before the impulse happens t=9.0− what is the total energy, in Joules, stored in the circuit? E. At the time just before the impulse happens what is the voltage vC(9.0−), in Volts, across the capacitor?

OpenStudy (anonymous):

How to calculate IL?

OpenStudy (anonymous):

4) iL (9) = cos(2*pi*f*t) For your example, frequency f = 0.25Hz so, iL = cos(2*pi*0.25*9)=0

OpenStudy (anonymous):

how to calculate vc?

OpenStudy (anonymous):

vc = -L*2*pi*f*sin(2*pi*f*t) Remember to keep your calculator mode in radian